2009-02-26 Telecon, 10:00AM EST (15:00 SAST)

Post-meeting notes:

-- Discussion of data highlights the following issues:

. Flatfields show large gradient (factor of 2 from center to edges), which is slightly worse since the mirror replacement (previous factor ~1.8)

. Flatfield pattern observed to rotates by 90. Consistent prior to 20081219, thenis "flopped". Datacube on 20090220 actually shows the flopoccurring over a few 10 sec. frames. Is this a telescope positional effect? Could an optic be loose and rotating?

. Optics are likely to be slightly misaligned due to mirror change. However, why is vignetting pattern still present?

. Ghosting was observed in recent data as well as engineering data. These are thought to be due to the dichroic. However, it is possilbe that new ghosts (that move opposite the images and are thus a different reflection) have appeared since engineering.

. There is a variable psf across the frame, which has been present since engineering. It is more apparent when there are bright sources.

-- Current image quality issues are serious and negatively impacting the data. We think that if we can sort out what is still causing the vignetting it will go a long way to solving the other problems.

Action Items:

(1) Amanda to determine location of spare lenses. Either these or a new set could be sent to MIT for optical bench testing of the lenses and path.

(2) Amanda and Mike to coordinate on Zemax modeling. Mike is taking a class in March; if he's well-prepared he can get even more help on our particular setup.

(3) Elisabeth & Carlos to continue data analysis and reporting issues that arise.

(4) Jim & Amanda Z. to prepare for observing run in March. Will avoid engineering, but might test new filter control software. To discuss beforehand what (if any) engineering data will fit into the plans

(5) Amanda to book travel to Hawaii in May (12-26). Need to contact Bobby/Lars concerning mounting brackets for filters (do they need to be machined to fit the new filters).

(6) Amanda to propose for Nix event 03 March.

 

MIT to call SA: +27.21.460.6294

participants: Gulbis (@ SAAO); Elliot, Person, Adams, Zuluaga, & Zangari (@ MIT)

Items to discuss:

(1) Current IRTF POETS issues
- flatfields (large gradient, variation with time, difficulty to obtain)
- psf variability across frame
- vignetting
- ghosting

(2) Addressing these problems
- complete characterization
- Zemax modeling
- onsite telescope and/or lab testing

(3) Filter control software update

(4) Onsite in Hawaii
- jle/amz visit in March (?)
- aasg visit in May (~12-26)

*Extra item: proposing for Nix 03 March event